# Deep learning reveals key predictors of thermal conductivity in covalent organic frameworks

### v1

@@ Abstract
The thermal conductivity of covalent organic frameworks (COFs), an emerging class of nanoporous polymeric materials, is crucial for many applications, yet the link between their structure and thermal properties is not well understood. From a dataset of over 2,400 COFs, we find that conventional features like density, pore size, void fraction, and surface area do not reliably predict thermal conductivity. To overcome this, we train an attention-based machine learning model that accurately predicts thermal conductivities, even for structures outside the training set. We then use the attention mechanism to understand why the model works. Surprisingly, dangling molecular branches emerge as key predictors of thermal conductivity, alongside conventional geometric descriptors like density and pore size. Our findings show that COFs with dangling functional groups exhibit lower thermal transfer capabilities than otherwise. Molecular dynamics simulations confirm this, revealing significant mismatches in the vibrational density of states due to the presence of dangling branches.
@@
Keywords
@@ Introduction
Covalent organic frameworks[1, 2] (COFs) are an emerging class of nanoporous polymeric materials. Compared to metal-organic framework (MOFs)[3, 4] and zeolites[5], the crystalline backbone of COFs is composed of organic building blocks – known as knots and linkers – connected by strong covalent bonds, thereby offering higher stability[6]. The crystalline nature of COFs, along with their high surface areas, tunable pore sizes, and functionalizable organic linkers, make them exceptionally suited for a wide range of promising applications. These applications include (but are not limited to) photoconductivity,[7, 8] chemo-sensing,[9, 10] catalysis,[11] drug delivery,[12, 13] thermoelectrics,[14, 15] semiconductors,[16] gas storage and separation.[17, 18]
@@
A key property dictating the applications of COFs is thermal transfer. For example, low thermal conductivity is desired for thermoelectrics to maintain large internal thermal gradients and increase efficiency.[19] On the other hand, a high thermal conductivity is desired for gas adsorption and separation where efficient thermal transfer is important for the longevity and stability of the nanoporous membranes.[20] COFs are promising nanoporous candidates for such applications as the ability to design their crystalline topology (e.g., lattice type, pore size) and chemistry via choice of knots and linkers opens up a large and diverse space of thermal conductivities.[21, 22, 23, 1, 2] Therefore, it is of great interest to gain a deeper understanding of the thermal transfer mechanism and to obtain strong structure-property trends, which in turn enables the application-specific design of COFs.
@@
Two approaches can be used to elucidate the thermal structure-property relationships of COFs: experimental trial-and-error methods which involve synthesis and characterization,[24, 25] and computational high-throughput screening which relies on first-principle calculations.[26, 27, 28] With COFs offering practically an unlimited design space, an experimental trial-and-error approach to screen new COF candidates (including developing new synthesis routes for each candidate) is prohibitively inefficient. Moreover, experimentally synthesized COFs will always contain crystalline defects,[29, 30] which greatly influence the thermal conductivity. This makes it difficult to relate the thermal transfer mechanisms to the geometrical and chemical make-up of a COF structure. On the contrary, molecular dynamics (MD) simulations [31] offer a high-throughput virtual screening alternative to lab-based experiments. However, despite recent advances in computing hardware, even virtual screening can be prohibitively inefficient for a large design space. This is highlighted by the example that generating a dataset of just 2,471 two-dimensional COFs and their thermal conductivity for this study required 1.3 million CPU-hours in cloud computing (translating into approximately four months and \$70,000 in time and cost, respectively). This underscores the need for efficient structure-property maps that are accurate, generalizable, and interpretable.
@@
Recently, Islamov et al.[32] performed a high-throughput screening study of over 10,000 MOF structures and their thermal conductivity ($\kappa$) using MD. They found that the majority of the MOFs possess $\kappa<1\ \text{W}\text{m}^{-1}\text{K}^{-1}$ with a few exceptions possessing ultra-high thermal conductivity ($\kappa>10\ \text{W}\text{m}^{-1}\text{K}^{-1}$). While COFs are generally more thermally stable (owing to the strong covalent bonding of the building blocks), a recent study by Thakur et al.[33] on high-throughput screening study on over 10,000 COFs demonstrated that, generally, COFs also possess a similar range of $\kappa$, with the majority of the structures exhibiting $\kappa<1\ \text{W}\text{m}^{-1}\text{K}^{-1}$. Interestingly, MOFs and COFs possess similar structure-property trends. (i) Increasing the pore size is mostly sufficient to achieve a low thermal conductivity. (ii) Additionally, one can increase the void fraction, the surface area, or include heavy atoms to further decrease thermal conductivity. (iii) However, to increase the thermal conductivity, there are many factors that need to align. Islamov et al.[32] demonstrated that while lowering the pore size and increasing the density results in a higher ceiling of attainable thermal conductivity, other factors, such as the topology, mass-mismatch, and linker length also need to be accounted for. Thakur et al.[33] came to similar conclusions, alongside the observation that aligning the polymeric chains in the structure to heat flow direction raises the ceiling of attainable thermal conductivity. However, these hand-crafted guidelines and ad hoc correlations are insufficient for developing an accurate and generalizable predictive model that captures the thermal conductivity structure-property relationships of COFs. Our study shows that no combination of commonly used descriptors consistently predicts thermal conductivity. Consequently, a definitive method for identifying COFs with tailor-made thermal properties for various applications remains elusive.
To address this knowledge gap, we turn to machine learning (ML) with an emphasis on interpretability and explainability. There are primarily two routes for ML modeling of structure-property maps of porous polymers (including but not limited to COFs) –
@@
  • high intepretability, limited accuracy: using hand-crafted and pre-extracted high-level of the crystalline network (e.g., pore size, density, surface area, atomic composition) as input to classical regression algorithms;[34, 35, 36]
@@
  • high accuracy, limited interpretability: using information-rich but raw graph representation of the crystalline network (where nodes denote atoms with features such as position and atom type, and edges denote bonds with the bond order as the feature) as input to end-to-end deep learning-based regression algorithms.[37, 38, 39, 40]
@@
In the following, we first conduct a large-scale data analysis of the thermal conductivity structure-property relations of COFs and identify the deficiencies in commonly used descriptors. Next, we provide a deep learning model that accurately predicts the thermal conductivity of COFs. Subsequently, an interpretability analysis of the deep learning model uncovers the presence of dangling atoms in the crystalline network of COFs as a strong and so-far missing key predictor for thermal conductivity. Additional physics-based analysis sheds light on the important role of dangling atoms in lowering the thermal conductivity of COFs by disrupting heat transfer pathways. We close by utilizing the ML model for efficient high-throughput screening and identifying COFs with extreme thermal conductivities.
@@ Results
We create a labeled dataset of COFs and their thermal conductivities by selecting 2,471 two-dimensional COFs from the unlabeled dataset of Mercado et al.[41] The COF structures in this design space are made up of one type of nodal linker (also referred to as knot) and one type of connecting linker. The sampled subset consists of COFs with 104 different linkers arranged in 25 different topologies (e.g., honeycomb, kagome, square lattice, etc.), and bonded by four different types of covalent linkages (i.e., carbon-carbon, amine, amide, and imine). We conduct non-equilibrium molecular dynamics (NEMD) simulations to calculate thermal conductivities in two orthogonal in-plane directions (see Supplementary Information Section 1.1 for more details). In Supplementary Information Section 1.2, we show that there is minimal in-plane anisotropy in thermal conductivity across the dataset. Therefore, we utilize average thermal conductivity in both directions, denoted as $\kappa$, as the quantity of interest to explore the structure-property relationships of COFs. In the following and in Figure 2, we present the observations from data analysis of the key descriptors of COFs and thermal conductivities in our labelled dataset.
@@
Figure 2: Distribution of (a) $\kappa$ versus density (with the color indicating data count per bin), (b) $\kappa$ versus largest pore diameter (LPD), (c) $\kappa$ versus void fraction, and (d) $\kappa$ versus gravimetric surface area (GSA). Also shown are four COF structures (i-iv) with contrasting properties, marked by green triangles in the plots above.
@@
Density: Figure 2a illustrates the distribution of $\kappa$ vs. density. There is a large spread of densities within the dataset and the majority of the COFs possess a density of 0.45-0.55 $\mathrm{g\ cm^{-3}}$. Moreover, in accordance with previous findings,[33] we observe an increasing trend in $\kappa$ with increasing density. However, increasing the density increases the ceiling of attainable $\kappa$, but does not guarantee a high value.
@@
Pore size: The pore size of any COF is represented by the largest pore diameter (LPD). Figure 2b shows that LPD in the dataset is spread over a large range between 5 Å and 80 Å and a concentration of structures in the 15-25 Å region. Moreover, there is an inverse relationship between the pore size and the thermal conductivity, as previously discovered by Freitas et al.,[42] where the range of achievable $\kappa$ decreases significantly with increasing pore sizes. Here again, a small pore size does not guarantee a high thermal conductivity. On the other hand, the pore size can be taken as the sole design parameter to effectively reduce the thermal conductivity range (as mentioned in refs.[33, 32]).
@@
Void fraction: Figure 2c shows that the distribution of the void fractions of COFs across the dataset are within the 0.5–0.96 range, with the biggest share of COFs possessing a void fraction of around 0.85 and higher. Furthermore, a high void fraction of around 0.85 delivers both the largest values and range of attainable $\kappa$ within our dataset. This finding is counterintuitive. Void fraction is typically inversely correlated with density (see Supplementary Information Section 1.3). Although low density is generally associated with a low $\kappa$ value, as shown in Figure 2a, the unexpected observation is that a high void fraction (which corresponds to low density) surprisingly results in a high $\kappa$ value, as also depicted in Figure 2c.
@@
Surface area: Gravimetric surface area (GSA), i.e., surface area per unit mass varies highly across all the COF structures (see Figure 2d) with a larger cluster of points in the 7000-8000 $\mathrm{m^{2}\ g^{-1}}$ range. We also observe that around intermediate GSA of 7,000 $\mathrm{m^{2}\ g^{-1}}$, we have the widest spread and highest values of $\kappa$.
@@
Poor $\kappa$-prediction with known descriptors: From the above initial analysis, we notice that the geometrical descriptors used here are insufficient to provide solid trends with respect to thermal conductivity. For example, with all the optimal descriptor values, i.e., a large density, a low pore size, a large void fraction, and an intermediate GSA, we are still not guaranteed to obtain a COF with a high $\kappa$. To elucidate this further, we choose four COFs – denoted by (i)-(iv) – and compare them in Figure 2a-d.
@@
The third COF (iii) has a pore size of 33.73 Å in the intermediate-to-high range, a low density at 0.408 $\mathrm{g\ cm^{-3}}$, a high void fraction at 0.932, and a slightly larger surface area of 8741 $\mathrm{m^{2}g^{-1}}$. Following the trends described in literature,[33, 42, 43] the geometrical descriptors are not optimal for a high $\kappa$. Nonetheless, COF (iii) has rather high $\kappa=1.96\ \mathrm{Wm^{-1}K^{-1}}$. Lastly, COF (iv) has an exceptionally high pore size at 80.68 Å and a low density of 0.262 $\mathrm{g\ cm^{-3}}$ but high void fraction of 0.958. Due to the extreme geometry, the COF (iv) has a very low thermal conductivity of $\kappa=0.289\ \mathrm{Wm^{-1}K^{-1}}$.
We further train four ensemble regression models as baselines based on Random Forest,[44] Gradient Boosting,[45] XGBoost,[46] and AdaBoost,[47] (see Supplementary Information Section 2 for training protocols and hyperparameters of each model) and features shown in Figure 3a to predict thermal conductivity. The same figure describes the corresponding goodness-of-fit $R^{2}$ for each model. The prediction accuracy is poor for all the models ($R^{2}<0.6$ across all models), and it is evident that these well-known descriptors are insufficient to predict the thermal conductivity. In summary, these examples illustrate the fact that the thermal conductivity structure-property relationships are highly complex and call for further examination.
@@
Figure 3: (a) Performance of ensemble regression models at predicting $\kappa$ using the shown descriptors. Average and standard goodness-of-fit $R^{2}$ are computed based on 10-fold cross-validation. (b) Schematic of PMTransformer model. A sample COF structure is shown on the left. A crystal graph convolutional neural network computes local embeddings of the COF graph, while the GRIDAY computes a three-dimensional energy grid of the structure, which becomes the global embeddings. These embeddings are combined and input into the PMTransformer (which includes a transformer encoder with attention mechanism and a prediction head) to predict $\kappa$. On the right, a parity plot compares $\kappa$ predictions with ground truth values. The dashed line represents the ideal line with zero intercept and unit slope.
@@
Our starting point is the deep learning-based Porous Material Transformer (PMTransformer) developed by Park et al.[40] The PMTransformer is a multi-modal transformer[48] model (see Figure 3b for schematic) designed for universal transfer learning for predicting the properties of porous materials, including COFs, MOFs, and zeolites. It receives two distinct sets of inputs: local and global features. Local features, which reflect the chemistry of building blocks and bonds, are derived from a crystal graph convolutional neural network[49] (CGCNN). The CGCNN directly processes the crystal graph, where atoms and bonds are nodes and edges, respectively, and generates embeddings that capture essential chemical details. These embeddings serve as local features in the PMTransformer. Conversely, global features describe crystalline characteristics, including topological and geometric descriptors such as pore size and surface area. These features are derived from three-dimensional energy grids, which are generated by calculating the interaction energy between the material structure and a methane gas molecule at each grid point using the package GRIDAY[50]. Similar to images in Vision Transformers[51], the energy grids are divided into patches and flattened through linear projections. Eventually, both local and global embeddings are input into a transformer encoder, where the attention mechanism[48] helps the model focus on the most relevant parts of the molecular structure for predicting (in our case) the thermal conductivity.
@@
To enable universal transfer learning, the transformer encoder in PMTransformer is pre-trained on an extremely large dataset of 1.9 million hypothetical porous materials to predict easily obtainable yet essential properties such as topology, void fraction, and building block prediction. This approach ensures the encoder captures critical information necessary for accurately predicting other, more complex properties in downstream tasks with much smaller datasets. For more details regarding the pre-training of PMTransfomer, refer to ref.[40]
@@
In this study, we leveraged the pre-trained transformer encoder of the PMTransformer model to perform transfer learning for the prediction of the thermal conductivity of COFs. We fine-tune the transformer encoder model using the pre-trained weights as a starting point and jointly train a prediction head based on a multi-layer perceptron architecture to map the output of the transformer encoder to the thermal conductivity of COFs. We used the mean squared error (MSE) as a loss function for training. The dataset of 2,471 COFs is randomly split into two subsets: 90% of the data is used for training (out of which 10% for validation), and the remaining data is used for testing. Additional details of the training protocol are provided in Supplementary Information Section 3.
We demonstrate the model’s prediction accuracy on the test dataset in Figure 3b. The model achieves a goodness-of-fit $R^{2}$ of $0.909\pm 0.006$ (across five random seeds) in predicting $\kappa$, with a mean absolute error (MAE) of $0.075\ \mathrm{Wm^{-1}K^{-1}}$. Compared to the four baseline regression models, which rely on only hand-crafted and/or commonly used features (as shown in Figure 3a), the deep learning model significantly outperforms them all. The accurate predictions on the test dataset not only highlight the model’s ability to capture the underlying structure-property relationships governing the thermal conductivity of COFs but also hints at the potential for uncovering the existing knowledge gap in those structure-property relationships.
@@
Additionally, we identify the main branch of a COF as the shortest continuous path connecting the boundary points necessary for periodicity. We then classify all the atoms extending and excluded from the main branch as dangling mass. More details on how we classify each atom as either part of the main branch or part of a dangling side branch are presented in Supplementary Information Section 4. We observe distinct patterns in the attention assigned to various atomic sites within a COF’s graph representation, which correspond to the location of the dangling masses. Two representative examples are shown in Figure 4.

### v2

@@ Abstract
The thermal conductivity of covalent organic frameworks (COFs), an emerging class of nanoporous polymeric materials, is crucial for many applications, yet the link between their structure and thermal properties remains poorly understood. Analysis of a dataset containing over 2,400 COFs reveals that conventional features such as density, pore size, void fraction, and surface area do not reliably predict thermal conductivity. To address this, an attention-based machine learning model was trained, accurately predicting thermal conductivities even for structures outside the training set. The attention mechanism was then utilized to investigate the model’s success. The analysis identified dangling molecular branches as a key predictor of thermal conductivity, a discovery supported by feature importance assessments conducted on regression models. These findings indicate that COFs with dangling functional groups exhibit lower thermal transfer capabilities. Molecular dynamics simulations support this observation, revealing significant mismatches in the vibrational density of states due to the presence of dangling branches.
@@
Covalent Organic Frameworks, Thermal Conductivity, Deep Learning, Attention
Prakash Thakolkaran1$\dagger$ Yiwen Zheng2$\dagger$ Yaqi Guo1 Aniruddh Vashisth2,* Siddhant Kumar1,*
{affiliations}
1Department of Materials Science and Engineering, Delft University of Technology, 2628 CD Delft, The Netherlands
2Department of Mechanical Engineering, University of Washington, Seattle, WA, USA
[…]
*Email: vashisth@uw.edu, sid.kumar@tudelft.nl; equal contribution
@@ Introduction
Covalent organic frameworks [1, 2] (COFs) are an emerging class of nanoporous polymeric materials. Compared to metal-organic framework (MOFs) [3, 4] and zeolites [5], the crystalline backbone of COFs is composed of organic building blocks – known as knots and linkers – connected by strong covalent bonds, thereby offering higher stability [6]. The crystalline nature of COFs, along with their high surface areas, tunable pore sizes, and functionalizable organic linkers, make them exceptionally suited for a wide range of promising applications. These applications include (but are not limited to) photoconductivity [7, 8], chemo-sensing [9, 10], catalysis [11], drug delivery [12, 13], thermoelectrics [14, 15], semiconductors [16], gas storage and separation [17, 18].
@@
A key property dictating the applications of COFs is thermal transfer. For example, low thermal conductivity is desired for thermoelectrics to maintain large internal thermal gradients and increase efficiency [19]. On the other hand, a high thermal conductivity is desired for gas adsorption and separation where efficient thermal transfer is important for the longevity and stability of the nanoporous membranes [20]. COFs are promising nanoporous candidates for such applications as the ability to design their crystalline topology (e.g., lattice type, pore size) and chemistry via choice of knots and linkers opens up a large and diverse space of thermal conductivities [21, 22, 23, 1, 2]. Therefore, it is of great interest to gain a deeper understanding of the thermal transfer mechanism and to obtain strong structure-property trends, which in turn enables the application-specific design of COFs.
@@
Two approaches can be used to elucidate the thermal structure-property relationships of COFs: experimental trial-and-error methods which involve synthesis and characterization [24, 25], and computational high-throughput screening which relies on first-principle calculations [26, 27, 28]. With COFs offering practically an unlimited design space, an experimental trial-and-error approach to screen new COF candidates (including developing new synthesis routes for each candidate) is prohibitively inefficient. Moreover, experimentally synthesized COFs will always contain crystalline defects [29, 30], which greatly influence the thermal conductivity. This makes it difficult to relate the thermal transfer mechanisms to the geometrical and chemical make-up of a COF structure. On the contrary, molecular dynamics (MD) simulations [31] offer a high-throughput virtual screening alternative to lab-based experiments. However, despite recent advances in computing hardware, even virtual screening can be prohibitively inefficient for a large design space. This is highlighted by the example that generating a dataset of just 2,471 two-dimensional COFs and their thermal conductivity for this study required 1.3 million CPU-hours in cloud computing (translating into approximately four months and \$70,000 in time and cost, respectively). This underscores the need for efficient structure-property maps that are accurate, generalizable, and interpretable.
@@
Recently, Islamov et al. [32] performed a high-throughput screening study of over 10,000 MOF structures and their thermal conductivity ($\kappa$) using MD. They found that the majority of the MOFs possess $\kappa<1\ \text{W}\text{m}^{-1}\text{K}^{-1}$ with a few exceptions possessing ultra-high thermal conductivity ($\kappa>10\ \text{W}\text{m}^{-1}\text{K}^{-1}$). While COFs are generally more thermally stable (owing to the strong covalent bonding of the building blocks), a recent study by Thakur et al. [33] on high-throughput screening study on over 10,000 COFs demonstrated that, generally, COFs also possess a similar range of $\kappa$, with the majority of the structures exhibiting $\kappa<1\ \text{W}\text{m}^{-1}\text{K}^{-1}$. Interestingly, MOFs and COFs possess similar structure-property trends. (i) Increasing the pore size is mostly sufficient to achieve a low thermal conductivity. (ii) Additionally, one can increase the void fraction, the surface area, or include heavy atoms to further decrease thermal conductivity. (iii) However, to increase the thermal conductivity, there are many factors that need to align. Islamov et al. [32] demonstrated that while lowering the pore size and increasing the density results in a higher ceiling of attainable thermal conductivity, other factors, such as the topology, mass-mismatch, and linker length also need to be accounted for. Thakur et al. [33] came to similar conclusions alongside the observation that aligning the polymeric chains in the structure to heat flow direction raises the ceiling of attainable thermal conductivity. However, these hand-crafted guidelines and ad hoc correlations are insufficient for developing an accurate and generalizable predictive model that captures the thermal conductivity structure-property relationships of COFs. Our study shows that no combination of commonly used descriptors consistently predicts thermal conductivity. Consequently, a definitive method for identifying COFs with tailor-made thermal properties for various applications remains elusive.
To address this knowledge gap, we turn to machine learning (ML) with an emphasis on interpretability and explainability. Previous efforts in ML-assisted design for thermally conductive organic materials include the discovery of polymers with high $\kappa$ through a hierarchical feature selection process [34], a reinforcement learning approach using SMILES-based representations [35], and methods leveraging molecular fingerprints as input features, such as fine-tuning a pre-trained regression model followed by screening [36], training a neural network [37], and employing an active learning approach [38]. Hu et al. [39] provide a review of recent efforts and outlook on ML for thermally conductive organic materials. For the ML modeling of structure-property maps of porous polymers (including but not limited to COFs) there are primarily two routes:
@@
  • high interpretability, limited accuracy: using hand-crafted and pre-extracted high-level of the crystalline network (e.g., pore size, density, surface area, atomic composition) as input to classical regression algorithms [40, 41, 42];
@@
  • high accuracy, limited interpretability: using information-rich but raw graph representation of the crystalline network (where nodes denote atoms with features such as position and atom type, and edges denote bonds with the bond order as the feature) as input to end-to-end deep learning-based regression algorithms [43, 44, 45, 46].
@@
In the following, we first conduct a large-scale data analysis of the thermal conductivity structure-property relations of COFs and identify the deficiencies in commonly used descriptors. Next, we provide a deep learning model that accurately predicts the thermal conductivity of COFs. Subsequently, an analysis of the attention scores of the deep learning model uncovers the presence of dangling atoms in the crystalline network of COFs as a strong and so-far missing key predictor for thermal conductivity. Additional physics-based analysis sheds light on the important role of dangling atoms in lowering the thermal conductivity of COFs by disrupting heat transfer pathways. We close by utilizing the ML model for efficient high-throughput screening and identifying COFs with extreme thermal conductivities.
@@ Results
We create a labelled dataset of COFs and their thermal conductivities by selecting 2,471 two-dimensional COFs from the unlabelled dataset of Mercado et al. [47] The COF structures in this design space are made up of one type of nodal linker (also referred to as knot) and one type of connecting linker. The sampled subset consists of COFs with 104 different linkers arranged in 25 different topologies (e.g., honeycomb, kagome, square lattice, etc.), and bonded by four different types of covalent linkages (i.e., carbon-carbon, amine, amide, and imine). We conduct non-equilibrium molecular dynamics (NEMD) simulations to calculate thermal conductivities in two orthogonal in-plane directions (see Supporting Information Section 1.1 for more details). In Supporting Information Section 1.2, we show that there is minimal in-plane anisotropy in thermal conductivity across the dataset. Therefore, we utilize average thermal conductivity in both directions, denoted as $\kappa$, as the quantity of interest to explore the structure-property relationships of COFs. In the following and in Figure 2, we present the observations from data analysis of the key descriptors of COFs and thermal conductivities in our labelled dataset. The distribution of all features and their correlations are detailed in Section 1.3 of the Supporting Information.
@@
Figure 2: Distribution of (a) $\kappa$ versus density (with the color indicating data count per bin), (b) $\kappa$ versus largest pore diameter (LPD), (c) $\kappa$ versus void fraction, and (d) $\kappa$ versus gravimetric surface area (GSA). Also shown are four COF structures (i-iv) with contrasting properties, marked by green triangles in the plots above. The $r$-value indicates the Pearson correlation coefficient.
@@
Density: Figure 2a illustrates the distribution of $\kappa$ vs. density. There is a large spread of densities within the dataset and the majority of the COFs possess a density of 0.45-0.55 $\mathrm{g\ cm^{-3}}$. Moreover, in accordance with previous findings [33], we observe an increasing trend in $\kappa$ with increasing density. The Pearson correlation coefficient between $\kappa$ and density is $r=0.594$, suggesting a moderate positive relationship. However, increasing the density increases the ceiling of attainable $\kappa$, but does not guarantee a high value.
@@
Pore size: The pore size of any COF is represented by the largest pore diameter (LPD). Figure 2b shows that LPD in the dataset is spread over a large range between 5 Å and 80 Å and a concentration of structures in the 15-25 Å region. Moreover, there is an inverse relationship between the pore size and the thermal conductivity, as previously discovered by Freitas et al. [48], where the range of achievable $\kappa$ decreases significantly with increasing pore sizes. This is evidenced by the correlation coefficient $r=-0.508$, indicating a moderate negative correlation. Here again, a small pore size does not guarantee a high thermal conductivity. On the other hand, the pore size can be taken as the sole design parameter to effectively reduce the thermal conductivity range (as mentioned in refs. [33, 32]).
@@
Void fraction: Figure 2c shows that the distribution of the void fractions of COFs across the dataset are within the 0.5–0.96 range, with the biggest share of COFs possessing a void fraction of around 0.85 and higher. Furthermore, a high void fraction of around 0.85 delivers both the largest values and range of attainable $\kappa$ within our dataset. This finding is counterintuitive. Void fraction is typically inversely correlated with density (see Supporting Information Section 1.3). Although low density is generally associated with a low $\kappa$ value, as shown in Figure 2a, the unexpected observation is that a high void fraction (which corresponds to low density) surprisingly results in a high $\kappa$ value, as also depicted in Figure 2c. Beyond a favorable range of void fractions, no distinct trend emerges in the relationship with $\kappa$, as the correlation coefficient $r=-0.213$ suggests only a weak negative relationship.
@@
Surface area: Gravimetric surface area (GSA), i.e., surface area per unit mass varies highly across all the COF structures (see Figure 2d) with a larger cluster of points in the 7000-8000 $\mathrm{m^{2}\ g^{-1}}$ range. We also observe that around intermediate GSA of 7,000 $\mathrm{m^{2}\ g^{-1}}$, we have the widest spread and highest values of $\kappa$. The correlation coefficient of $r=-0.147$ indicates a very weak negative relationship, suggesting that GSA is not a significant predictor.
@@
From the above initial analysis, we notice that the geometrical descriptors used here are insufficient to provide solid trends with respect to thermal conductivity. For example, with all the optimal descriptor values, i.e., a large density, a low pore size, a large void fraction, and an intermediate GSA, we are still not guaranteed to obtain a COF with a high $\kappa$. To elucidate this further, we choose four COFs – denoted by (i)-(iv) – and compare them in Figure 2a-d.
@@
The third COF (iii) has a pore size of 33.73 Å in the intermediate-to-high range, a low density at 0.408 $\mathrm{g\ cm^{-3}}$, a high void fraction at 0.932, and a slightly larger surface area of 8741 $\mathrm{m^{2}g^{-1}}$. Following the trends described in literature [33, 48, 49], the geometrical descriptors are not optimal for a high $\kappa$. Nonetheless, COF (iii) has rather high $\kappa=1.96\ \mathrm{Wm^{-1}K^{-1}}$. Lastly, COF (iv) has an exceptionally high pore size at 80.68 Å and a low density of 0.262 $\mathrm{g\ cm^{-3}}$ but high void fraction of 0.958. Due to the extreme geometry, the COF (iv) has a very low thermal conductivity of $\kappa=0.289\ \mathrm{Wm^{-1}K^{-1}}$.
Furthermore, we trained classical ensemble regression models using these descriptors to predict thermal conductivity (detailed results provided in Supporting Information Section 2) and observed that the models yield poor prediction accuracy. This result highlights the limitations of these descriptors in capturing the structure-property relationships underlying thermal conductivity. In summary, the correlation analysis with the previously introduced descriptors and the presented examples illustrate the fact that the thermal conductivity structure-property relationships are complex and call for further examination.
@@
Figure 3: Schematic of PMTransformer model. A sample COF structure is shown on the left. A crystal graph convolutional neural network computes local embeddings of the COF graph, while GRIDAY computes a three-dimensional energy grid of the structure, which becomes the global embeddings. These embeddings are combined and input into the PMTransformer (which includes a transformer encoder with attention mechanism and a prediction head) to predict $\kappa$. On the right, a parity plot compares $\kappa$ predictions with ground truth values. The dashed line represents the ideal line with zero intercept and unit slope.
@@
Our starting point is the deep learning-based Porous Material Transformer (PMTransformer) developed by Park et al. [46] The PMTransformer is a multi-modal transformer [50] model (see Figure 3 for schematic) designed for universal transfer learning for predicting the properties of porous materials, including COFs, MOFs, and zeolites. It receives two distinct sets of inputs: local and global features. Local features, which reflect the chemistry of building blocks and bonds, are derived from a crystal graph convolutional neural network [51] (CGCNN). The CGCNN directly processes the crystal graph, where atoms and bonds are nodes and edges, respectively, and generates embeddings that capture essential chemical details. These embeddings serve as local features in the PMTransformer. Conversely, global features describe crystalline characteristics, including topological and geometric descriptors such as pore size and surface area. These features are derived from three-dimensional energy grids, which are generated by calculating the interaction energy between the material structure and a methane gas molecule at each grid point using the package GRIDAY [52]. Similar to images in Vision Transformers [53], the energy grids are divided into patches and flattened through linear projections. Eventually, both local and global embeddings are input into a transformer encoder, where the attention mechanism [50] helps the model focus on the most relevant parts of the molecular structure for predicting (in our case) the thermal conductivity.
@@
To enable universal transfer learning, the transformer encoder in PMTransformer is pre-trained on an extremely large dataset of 1.9 million hypothetical porous materials to predict easily obtainable yet essential properties such as topology, void fraction, and building block prediction. This approach ensures the encoder captures critical information necessary for accurately predicting other, more complex properties in downstream tasks with much smaller datasets. For more details regarding the pre-training of PMTransfomer, refer to [46].
@@
In this study, we leveraged the pre-trained transformer encoder of the PMTransformer model to perform transfer learning for the prediction of the thermal conductivity of COFs. We fine-tune the transformer encoder model using the pre-trained weights as a starting point and jointly train a prediction head based on a multi-layer perceptron architecture to map the output of the transformer encoder to the thermal conductivity of COFs. We used the mean squared error (MSE) as a loss function for training. The dataset of 2,471 COFs is randomly split into two subsets: 90% of the data is used for training (out of which 10% for validation), and the remaining data is used for testing. Additional details of the training protocol are provided in Supporting Information Section 3.
We demonstrate the model’s prediction accuracy on the test dataset in Figure 3, where it achieves a goodness-of-fit $R^{2}$ of $0.909\pm 0.006$ (across five random seeds) and a mean absolute error (MAE) of $0.075\ \mathrm{Wm^{-1}K^{-1}}$ for predicting $\kappa$. The model’s strong performance suggest that there may be gaps in the feature sets typically used to describe COFs. It also highlights the possibility that additional key descriptors, beyond those commonly used, could play an important role in predicting thermal conductivity, thus warranting further exploration of structure-property relationships.
@@
Additionally, we identify the main branch of a COF as the shortest continuous path connecting the boundary points necessary for periodicity. We then classify all the atoms extending and excluded from the main branch as dangling mass. More details on how we classify each atom as either part of the main branch or part of a dangling side branch are presented in Supporting Information Section 4. We observe distinct patterns in the attention assigned to various atomic sites within a COF’s graph representation, which correspond to the location of the dangling masses. Two representative examples are shown in Figure 4.